OpenStudy (kdvldhvlaid):

Solve for x: −4|x + 5| = −16

OpenStudy (torialaw192):

Okay, what do you get if you subtract 5 from both sides?

jimthompson5910 (jim_thompson5910):

You have to divide both sides by -4 first

OpenStudy (legomyego180):

yup

OpenStudy (radar):

Is this a complex numbers problem? is the I actually i (for imaginary)

OpenStudy (legomyego180):

how could it be complex? there is no radical

OpenStudy (legomyego180):

Oh I see, no its an absolute value problem, @radar

OpenStudy (legomyego180):

Since nobody else is explaining it ill have a go \[|x+5|=4\] Recall, numbers inside of the absolute value always have to be positive, this is the fact that will allow us to solve this problem. The obvious answer is: x= -1 here right? Plug in -1 for x you get |-1+5| = 4 which satisfies the equation. The not so obvious thing we need to realize is that |-4| = 4 So in addition to solving for 4 we also need to solve for -4. |-9+5| = |-4| = 4 So x =-1, x =-9

jimthompson5910 (jim_thompson5910):

Another method is to break up |x+5| = 4 into the two equations x+5 = 4 x+5 = -4 then solving each for x. You'll get the same two answers. The rule I'm using is `if |x| = k then x = k or x = -k where k is some positive number` eg: if |x| = 5, then x = 5 or x = -5

OpenStudy (radar):

Sorry if brought confusion to the problem. I mistook the I for a capital i, and overlooked that it was an absolute value problem.

OpenStudy (mathmale):

@kdvldhvlaid: Hope this input has been helpful. Please at least acknowledge others' efforts to help you, and please close your post when you're satisfied with your result. Thank you.